OpenStudy (caitlinnr14):

Solve y/-2 -3 = 12 30 −30 18 −18

OpenStudy (anonymous):

To solve this question, first move -3 to the other side of the equation. Hence becoming this. y/-2=12+3 -> y/-2=15 After doing this, multiply -2 on both sides in order to get rid of -2 on the denominator that would give you y=-30

OpenStudy (caitlinnr14):

wait i got 2 different answers @varun1495 @SAValkyrie

OpenStudy (anonymous):

Varun y+6=-24 , when you solve that you would get y=-30

OpenStudy (anonymous):

on doing lcm we get (y+6)/-2=12 y+6=-24 y=-30

OpenStudy (caitlinnr14):

oki thx